A skeleton has a carbon-14 decay rate of o.27 disintegrations per minute per a gram of carbon. When did the animal live Assume that living organisms have carbon-14 decay rate of 15.3 d/min g Cand that carbon-14 has a half-life of 5730 yrs.
Answer #1

Apply half life equation

A = A0*(1/2)^(t/HL)

0.27 = 15.3*(0.5)^(t/5730)

t = 5730* ln(0.27 /15.3 ) / ln(0.5)

t = 33373.974 years old